Please help me with this problem. After doing the problem my answers below were wrong. Show all work and steps. Thanks

Maricopa's Success scholarship fund receives a gift of $ 170000. The money is invested in stocks, bonds, and CDs. CDs pay 3.25 % interest, bonds pay 2.2 % interest, and stocks pay 7.3 % interest. Maricopa Success invests $ 10000 more in bonds than in CDs. If the annual income from the investments is $ 7325 , how much was invested in each account?
Maricopa Success invested $ 63500 in stocks.
Maricopa Success invested $ 73500 in bonds.
Maricopa Success invested $ 33000 in CDs.




Answer by Boreal(15235)

x=bonds
stocks+CDs=170000-x
but x-10000=CDs
therefore, stocks+(x-10000)=170000-x
stocks=180000-2x
0.022x+0.0325(x-10000)+0.073(180000-2x)=170000
0.022x-325+0.0325x+13140-0.146x=7325
-0.0915x=-5490
x=$60,000 in bonds earn 1320
x-10000=$50,000 in CDs, earn 1625
180000-2x=$60,000 in stocks earn 4380. They add to $7325.